Reports 1-1 of 1 Clear search Modify search
MIF (General)
takaaki.yokozawa - 7:43 Friday 04 April 2025 (33232) Print this report
Alignment work 250404
After the large earthquake, the alignment became worse and LSC_LOCK guardian stopped at the PRMI lock trial with ALS_DARM,CARM lock.
So, I performed the initial alignment only for the PRMI, then PRFPMI_RF_LOCKED succeeded.
Next, I tried to perform the OMC alignment with the following klog
https://klog.icrr.u-tokyo.ac.jp/osl/?r=33214
Then, interferometer became stable.

Today, IPC glitch like signal detected so much (once per several minutes).
Search Help
×

Warning

×